Mary would like to save $10 000 at the end of 5 years for a future down payment on a car.

How much should she deposit at the end of each week in a savings account that pays 1.2%/a, compounded monthly, to meet her goal?
1
Expert's answer
2019-10-02T10:02:03-0400

Define by the formula:

S =D [ (1+r)^n -1 ] /r

S - what need at the end, 10,000

D - the monthly deposit (divide it by four for weekly)

n - the number of compounding periods, 5*12 = 60

r - the interest rate (decimal) at each compounding period, 0.12/12 = 0.01

so

10,000 = D[ 1.01)^60 -1 ] / 0.01

10,000 = D (0.817)/0.01 = D (81.7)

D = $122.44 per month

D/4 = $30.61 per week